### v3

# CodeNER: Code Prompting for Named Entity Recognition

Abstract
Recent studies have explored various approaches for treating candidate named entity spans as both source and target sequences in named entity recognition (NER) by leveraging large language models (LLMs). Although previous approaches have successfully generated candidate named entity spans with suitable labels, they rely solely on input context information when using LLMs, particularly, ChatGPT. However, NER inherently requires capturing detailed labeling requirements with input context information. To address this issue, we propose a novel method that leverages code-based prompting to improve the capabilities of LLMs in understanding and performing NER. By embedding code within prompts, we provide detailed BIO schema instructions for labeling, thereby exploiting the ability of LLMs to comprehend long-range scopes in programming languages. Experimental results demonstrate that the proposed code-based prompting method outperforms conventional text-based prompting on ten benchmarks across English, Arabic, Finnish, Danish, and German datasets, indicating the effectiveness of explicitly structuring NER instructions. We also verify that combining the proposed code-based prompting method with the chain-of-thought prompting further improves performance. Corresponding author.Our code and dataset are available at https://github.com/HanSungwoo/CodeNER.
CodeNER: Code Prompting for Named Entity Recognition
Sungwoo Han1, Jingun Kwon1, Hidetaka Kamigaito2, and Manabu Okumura3 1Chungnam National University, 2Nara Institute of Science and Technology (NAIST) 3Institute of Science Tokyo 77sungwhan@o.cnu.ac.kr jingun.kwon@cnu.ac.kr kamigaito.h@is.naist.jp oku@pi.titech.ac.jp
1 Introduction
Named entity recognition (NER) is a task of detecting and classifying named entities in a given text, such as person, location, and miscellaneous entities. Traditionally, NER has been treated as a sequential labeling problem, e.g., with the BIO output schema, where a learned model classifies each token in a sequence from a predefined label set carreras-etal-2003-learning; lample-etal-2016-neural; ma-hovy-2016-end; wan-etal-2022-nested. These schemata indicate that each token in the input should be aligned with the BIO tag by capturing entity boundaries. Incorporating conditional random field (CRF) networks, which explicitly capture broader context information of input, has led to leading performance 10.5555/645530.655813; finkel-etal-2005-incorporating; liu-etal-2011-recognizing; lample-etal-2016-neural; jie-lu-2019-dependency; kruengkrai-etal-2020-improving; xu-etal-2021-better.
Recently, large language models (LLMs) have achieved great success in various natural language processing (NLP) tasks by leveraging in-context learning with prompts and exploiting their zero-shot task-solving abilities without the need to update model parameters neuips2022; weifinetuned; 51115. There are two main types of LLMs: closed and open. While closed models, such as ChatGPT, which is based on GPT-4, [^1] https://chat.openai.com/ have shown strong reasoning abilities, open models, such as Llama-3 and Phi-3, offer flexibility to access and modify their architectures llama3modelcard; abdin2024phi3technicalreporthighly.
Motivated by these recent advancements, the zero-shot reasoning abilities of LLMs, specifically GPT, in NER have been explored xie-etal-2023-empirical; 10.1109/JCDL57899.2023.00034; xie-etal-2024-self; 10.1007/978-3-031-72437-4_22. Despite the success of using LLMs, however, previous work has relied on two-stage approaches, such as majority voting or constructing self-annotated datasets, for few-shot techniques with text-based prompting. Text-based prompting faces challenges in NER, primarily due to the inherent gap between the text-in-text-out schema of LLMs and the text-in-span-out nature of the task 10.5555/AAI28114605. This mismatch makes it difficult to handle the sequential aspect of the BIO output schema, which is integral to NER, in LLMs and can result in poor performance in identifying entity boundaries ding-etal-2024-rethinking. Although various prompting methods, such as chain-of-thought (CoT) han2023informationextractionsolvedchatgpt; xie-etal-2023-empirical, have been introduced to improve reasoning cot, challenges still remain in bridging this gap.
In this paper, we propose CodeNER, a code-based prompting method designed to improve the capabilities of LLMs in understanding and performing NER. While previous supervised frameworks learn an explicit mapping from each token to its label, this mapping is not captured in zero-shot and few-shot settings, specifically in-context learning. By embedding detailed BIO schema labeling instructions within complete, structured code prompts, CodeNER better identifies entity boundaries by exploiting the LLMs’ inherent ability to comprehend programming language structures. Unlike previous text-based or partial code-based approaches, our method explicitly guides the sequential processing essential for accurate NER, thereby bridging the gap between the text-in-text-out schema of LLMs and the text-in-span-out nature of the task. This explicit sequential guidance is particularly crucial in zero-shot and few-shot scenarios, where explicit supervision is minimal. Completely structured code prompts with BIO out schema inherently capture necessary sequential context, addressing the limitations of zero-shot and few-shot methods, which typically lack explicit sequential modeling due to the absence of direct token-label supervision.
We evaluate the performance of CodeNER compared with the text-based prompting methods for both closed models of ChatGPT (GPT-4 and GPT-4 Turbo) and open models of Llama-3-8B and Phi-3-mini-128k-instruct. Experimental results obtained on ten benchmarks across English, Arabic, Finnish, Danish, and German NER datasets demonstrated that the proposed CodeNER outperforms the text-based prompting methods. We perform in-depth analysis to evaluate the effectiveness of CodeNER in terms of overcoming the limitations of the text-based prompting methods, that rely solely on context information of input in zero-shot settings. We also show that the performance of the proposed method can be further improved by incorporating CoT into CodeNER.
2 Related Work
Named Entity Recognition. NER is the task of identifying and annotating named entities in a given text from a predefined label set. Traditionally, NER has been approached as a sequential labeling task carreras-etal-2003-learning; lample-etal-2016-neural. Early efforts employed LSTM networks to represent token embeddings and subsequently integrated conditional random field (CRF) layers with CNNs to consider broader context information collobert. In addition, hybrid models that combine LSTM and CNNs have been investigated to leverage character-level information chiu-nichols-2016-named; ma-hovy-2016-end. Another approach formulates the NER task as span prediction, where learned models simultaneously detect span boundaries and annotate named entities stratos-2017-entity; li-etal-2020-unified; ouchi-etal-2020-instance; fu-etal-2021-spanner. Recent work has extended the span-based approaches to better capture nested entities, which can appear when one named entity is contained within another shen-etal-2021-locate; tan2021sequencetoset; ding-etal-2022-ask. Instead of formulating NER as a sequential labeling task, generative methods have been introduced, that utilize a Seq2Seq framework by considering candidate named entity spans as both source and target sequences cui-etal-2021-template; Fei_Ji_Li_Liu_Ren_Li_2021; yan-etal-2021-unified-generative. While supervised methods have achieved saturated performance on the standard NER benchmarks, they require extensive labeled datasets chen-etal-2021-data; zhou-etal-2022-melm and domain adaptation methods zhang-etal-2021-pdaln; yang-etal-2022-factmix; xu-etal-2023-improving.
In-context Learning. LLMs have achieved significant success in various NLP tasks in zero-shot settings by leveraging specifically formatted instruction-based inputs Radford2019LanguageMA; NEURIPS2020_1457c0d6; chowdhery2022palm. Different from the fine-tuning setting, which requires additional parameter updates for learning to solve tasks, the zero-shot setting leverages in-context learning to perform tasks without any additional adjustments of the parameters of a pre-trained model. Owing to its flexibility, numerous efforts have been made to investigate better instructions for enhancing reasoning abilities, including few-shot, directional stimulus, and CoT learning liu-etal-2022-generated; zhu2023multilingual; cot; li2023guiding. Recently, the use of LLMs in NER has gained attention due to their remarkable reasoning abilities. wang2023gptnernamedentityrecognition considered a retrieval-augmented generation approach by leveraging entity-level information using a fine-tuned NER model to search exemplars. xie-etal-2023-empirical explored the reasoning abilities of ChatGPT by considering few-shot techniques based on a two-stage majority voting. xie-etal-2024-self demonstrated that self-annotating an unlabeled dataset using ChatGPT further improves performance with the two-stage majority voting. Additionally, instruction-based fine-tuning with LLMs and evaluating unseen labels on new datasets using different prompts have also been studied sainz2024gollie; ding-etal-2024-rethinking; zamai2024lessinstructmoreenriching.
Figure 1: Examples of text- and code-based prompts for zero-shot NER with ChatGPT. (a) and (b) indicate the text-based and code-based prompts, respectively.
In this study, instead of leveraging a text-based approach, we focus on leveraging a programming language-style for NER to exploit the structured nature inherent in code. Previous studies have primarily focused on broader context information of input and have not thoroughly explored the sequential aspect of BIO output schema into NER prompting. The proposed CodeNER bridges this gap by integrating the BIO output schema and broader context of input in LLMs to improve NER performance.
3 CodeNER
Unlike text-based prompting methods, which rely solely on natural language instructions for NER, our CodeNER is a code-based prompting method. Figures 1 (a) and (b) show text-based prompts and code-based prompts in Python, respectively. CodeNER explicitly encourages models to consider the BIO output schema and broader context information of input via prompts written in a programming language-style. This approach is critical for improving performance in NER tasks since it enables the model to better capture entity boundaries ding-etal-2024-rethinking than existing methods using LLMs that consider only context information liu-etal-2011-recognizing; kruengkrai-etal-2020-improving; xu-etal-2021-better. Python code-based prompts provide a clear and human-readable structure while overcoming several challenges, such as potential misinterpretation and variability in natural languages wang-etal-2023-code4struct; LiSTYWHQ23; sainz2024gollieannotationguidelinesimprove.
In CodeNER, the variable “sentence” contains a list of sentences tokenized for processing, and “ner_tag_labels” is a dictionary that maps each NER label to a numeric identifier with comments that explain each label. The prompt includes a function that performs NER, followed by a print statement that outputs the annotated named entities. The function “find_ner_tags” uses a for-loop to guide the LLM to process the BIO output schema and input sentence information. [^2] We performed preliminary tests and found that the function is indeed essential for improving performance and facilitating accurate parsing. In addition, “ner_word_dictionary,” which is not predefined, is included with explanatory comments to prompt the model to define it dynamically for NER. The classified results are appended to the list “ner_tags_dict_tags,” and the final output is generated.
By embedding the prompt within a code structure, CodeNER not only specifies the BIO output schema but also incorporates input sentence information, providing a comprehensive framework for the model to perform NER.
4 Experiments
4.1 Experimental Settings
Language Dataset #Sent. #Label Language Dataset #Sent. #Label
English CoNLL03 3,453 5 Arabic Arabic A 267 5
FIN 305 5 Arabic B 254 5
MIT Movie Trivia 1,953 13 Finnish finnish A 986 5
MIT restaurant 1,521 9 Danish DaNE 565 5
WNUT 17 1,287 7 German SwissNER 200 5
Table 1: Statistics of the datasets. #Sent. and #label indicate the dataset size and the number of named entities, respectively.
Datasets. We used ten benchmark datasets. Table 1 provides the statistics of the NER test datasets. For the English datasets, we used CoNLL03 tjong-kim-sang-de-meulder-2003-introduction, FIN salinas-alvarado-etal-2015-domain, WNUT-17 derczynski-etal-2017-results, the MIT movie review semantic corpus, and the MIT restaurant review. [^3] The MIT NER datasets are publicly available at https://groups.csail.mit.edu/sls/downloads/. For the Arabic datasets, we used ArabicDatasets A and B bari19. For the Finnish dataset, we used FinnishDataset A Ruokolainen_2019. The German and Danish datasets are SwissNER vamvas-etal-2023-swissbert and DaNE hvingelby-etal-2020-dane, respectively.
We transformed the traditional NER datasets into a programming language-style format. Note that, when considering languages other than English, we used Google Translate to translate the “ner_tags_labels” descriptions in both CodeNER and Vanilla prompts. [^4] https://translate.google.com/ Different from previous work that sampled only a small portion of datasets xie-etal-2023-empirical; xie-etal-2024-self, we used the entire datasets to ensure a comprehensive evaluation.
Implementation Details. For closed models, we employed the recent ChatGPT (GPT-4 and GPT-4 Turbo) models with their official APIs. For open models, we employed open-source Llama-3-8B and Phi-3-mini-128k-instruct as the backbone models for our experiments. Following previous work, we used F1 scores across all datasets based on the model-generated outputs cui-etal-2021-template; xie-etal-2023-empirical by considering the traditional evaluation method of BIO tagging hammerton-2003-named; ma-hovy-2016-end; chen-etal-2020-local. All experiments for the open models were conducted using an NVIDIA RTX A6000 GPU. For the decoding step, we set the $\rm do_{sample}$ option to False. [^5] All hyperparameter settings are described in Appendix A.
Baseline. We compared the proposed CodeNER to vanilla text-based prompting NER, shown in Figure 1 (a) xie-etal-2023-empirical in zero-shot settings. [^6] We converted model-generated outputs to the BIO format for evaluation.
4.2 Main Results
GPT-4 GPT-4 turbo
Language Dataset Vanilla CodeNER $\Delta$ Vanilla CodeNER $\Delta$
English CoNLL03 73.24 69.48 -3.76 74.54 68.31 -6.23
FIN 22.83 41.13 18.30 11.58 27.45 15.87
MIT restaurant 40.47 49.95 9.48 36.34 46.05 9.71
MIT movie 52.21 48.38 -3.83 50.78 46.80 -3.98
WNUT-17 51.04 46.00 -5.04 42.80 52.09 9.29
Arabic Arabic A 38.31 44.52 6.21 39.11 41.58 2.47
Arabic B 43.04 45.77 2.73 44.29 46.57 2.28
Finnish Finnish A 54.26 59.85 5.59 49.32 62.98 13.66
German SwissNER 58.18 61.74 3.56 66.71 57.22 -9.49
Danish DaNE 66.46 57.28 -9.18 65.53 62.61 -2.92
Average 50.00 52.41 2.41 48.10 51.17 3.07
Table 2: Experimental results with macro-F1 scores on ten benchmarks. $\Delta$ indicates the performance difference between CodeNER and Vanilla.
PER LOC ORG MISC
Dataset Model Vanilla CodeNER $\Delta$ Vanilla CodeNER $\Delta$ Vanilla CodeNER $\Delta$ Vanilla CodeNER $\Delta$
CoNLL03 GPT-4 92.06 88.38 -3.68 79.45 75.77 -3.68 66.15 58.53 -7.62 31.88 36.87 4.99
GPT-4 turbo 90.14 86.16 -3.98 83.19 72.19 -11.00 63.96 55.20 -8.76 43.07 48.99 5.92
FIN GPT-4 26.94 51.71 24.77 15.62 38.10 22.48 14.72 7.48 -7.24 1.06 0.78 -0.28
GPT-4 turbo 5.29 31.76 26.47 30.77 27.72 -3.05 23.93 13.98 -9.95 0.00 0.79 0.79
Arabic A GPT-4 63.20 71.10 7.90 52.99 55.64 2.65 21.38 15.31 -6.07 16.91 26.53 9.62
GPT-4 turbo 54.08 68.33 14.25 59.06 57.30 -1.76 27.37 18.39 -8.98 18.90 19.42 0.52
Arabic B GPT-4 63.97 67.94 3.97 57.43 59.73 2.30 21.62 21.05 -0.57 23.66 26.42 2.76
GPT-4 turbo 62.70 68.47 5.77 61.68 64.18 2.50 27.27 26.75 -0.52 24.07 24.26 0.19
Finnish A GPT-4 73.57 62.48 -11.09 63.81 81.83 18.02 60.20 66.87 6.67 12.83 14.73 1.90
GPT-4 turbo 65.75 82.32 16.57 62.88 79.84 16.96 53.78 68.06 14.28 9.02 14.92 5.90
SwissNER GPT-4 65.00 72.66 7.66 75.26 72.30 -2.96 60.75 63.23 2.48 20.00 32.16 12.16
GPT-4 turbo 83.00 74.07 -8.93 80.00 71.39 -8.61 73.96 52.36 -21.60 17.93 27.18 9.25
DaNE GPT-4 85.13 63.67 -21.46 76.15 66.12 -10.03 70.49 60.70 -9.79 25.66 36.20 10.54
GPT-4 turbo 83.91 79.77 -4.14 85.56 69.23 -16.33 67.91 63.19 -4.72 19.12 31.06 11.94
Average GPT-4 67.12 68.28 1.16 60.10 64.21 4.11 45.04 41.88 -3.16 18.86 24.81 5.95
GPT-4 turbo 63.55 70.13 6.58 66.16 63.12 -3.04 48.31 42.56 -5.75 18.87 23.80 4.93
Table 3: Experimental results with F1-scores for each label on CoNLL03, FIN, ArabicDatasetA, ArabicDatasetB, finnishDatasetA, DaNE, and SwissNER, where the label set includes PER, LOC, ORG, and MISC.
Table 2 shows the zero-shot results of CodeNER and Vanilla using GPT-4 and GPT-4 Turbo. For GPT-4, CodeNER outperformed the Vanilla method on most datasets, particularly with a significant improvement on FIN, where CodeNER achieved an 18.3 point improvement in the $F_{1}$ score. Similarly, a significant improvement was obtained on MIT restaurant, with an increase of 9.48 points. However, CodeNER underperformed GPT-4 on CoNLL03 and WNUT-17. On average, CodeNER improved the $F_{1}$ score by 2.41 points, representing a 4.82% improvement across all datasets. For GPT-4 turbo, similar performance gains were obtained by CodeNER, with an average improvement of 3.07 points in the $F_{1}$ score, representing a 6.37% improvement across all datasets.
Table 3 shows a detailed comparison between CodeNER and Vanilla across four NER labels: PER (Person), LOC (Location), ORG (Organization), and MISC (Miscellaneous) on CoNLL03, FIN, Arabic A, Arabic B, Finnish A, DaNE, and SwissNER. On average, the proposed CodeNER generally outperformed Vanilla, except for the ORG label. In particular, we observed that CodeNER consistently outperforms Vanilla for the MISC label. Although there exist labels for which CodeNER underforms ChatGPT on certain datasets, specifically CoNLL03 and SwissNER, the overall performance demonstrated that CodeNER, which can incorporate the BIO output schema together, is more robust for NER in zero-shot settings. Table 9 shows the additional results for one-shot and few-shot settings.
Figure 2 shows a detailed comparison for each label between CodeNER and Vanilla on MIT-restaurant and MIT-movie. On MIT-restaurant, CodeNER consistently outperformed Vanilla across most of the labels. However, on MIT-movie, Vanilla outperformed CodeNER for the “Plot” and “Quote” labels even though CodeNER maintained competitive scores.
Figure 2: Experimental results for each label on MIT-Restaurant and MIT-Movie.
Model Dataset Template $F_{1}$ PER LOC ORG MISC Model $F_{1}$ PER LOC ORG MISC Model $F_{1}$ PER LOC ORG MISC
Phi-3 FIN Vanilla 1.24 1.82 0.00 0.00 0.00 Llama-3 8B 18.14 16.47 35.44 15.00 0.00 Llama-3 70B 14.85 9.96 40.91 17.26 4.55
GoLLIE 8.26 9.59 6.90 5.02 0.00 9.21 9.87 8.60 8.22 0.00 11.02 6.84 36.84 10.96 0.00
GNER 26.98 30.50 27.35 16.36 0.44 33.18 44.14 14.58 7.38 0.00 24.90 29.05 27.35 10.00 1.23
CodeNER 17.51 18.18 27.12 10.43 0.00 31.96 39.22 34.04 6.54 0.00 19.39 18.46 37.25 13.02 1.13
CodeNER w/o label 8.49 7.35 19.05 6.62 0.00 45.62 57.26 41.12 9.40 1.33 26.11 29.29 36.17 9.89 1.49
CoNLL03 Vanilla 11.73 21.20 15.44 3.75 0.00 64.02 88.90 71.01 54.85 11.85 68.13 93.39 74.01 57.40 21.32
GoLLIE 35.39 51.47 33.88 35.43 1.84 34.72 49.03 38.32 30.65 2.84 53.93 72.04 60.05 44.54 19.87
GNER 44.63 69.87 50.31 31.88 3.14 34.43 44.81 49.12 49.12 1.06 53.38 77.48 58.52 38.96 19.79
CodeNER 36.57 54.03 44.64 19.75 16.95 40.29 51.07 50.70 31.01 12.71 54.94 76.42 57.48 43.45 26.61
CodeNER w/o label 48.33 66.92 55.66 32.26 26.12 37.00 44.27 46.17 31.80 10.75 47.53 60.61 53.89 36.43 28.59
SwissNER Vanilla 4.14 5.45 3.23 5.71 1.72 49.04 72.91 58.82 52.06 6.15 59.82 80.18 71.09 66.30 11.85
GoLLIE 32.54 51.64 40.51 32.38 2.42 24.98 38.58 29.10 24.56 6.51 44.49 54.78 62.54 42.62 9.15
GNER 22.37 31.73 26.84 21.36 8.29 25.58 28.27 38.34 25.74 2.29 43.07 55.90 58.67 40.10 11.21
CodeNER 28.76 49.47 42.34 14.35 12.44 35.48 28.99 62.37 30.55 6.55 47.04 66.92 61.57 42.23 13.55
CodeNER w/o label 31.78 50.55 41.67 21.71 15.76 31.87 25.37 54.41 28.09 8.02 39.21 58.52 52.50 31.90 12.50
Arabic A Vanilla 0.00 0.00 0.00 0.00 0.00 8.88 6.13 15.53 6.25 6.31 27.25 49.61 34.62 23.02 10.24
GoLLIE 14.60 16.74 22.80 8.30 8.76 12.57 26.4 16.36 7.25 2.99 22.83 32.84 35.46 14.14 9.89
GNER 4.23 3.61 1.75 0.00 6.92 6.53 12.09 11.93 0.00 0.71 22.66 37.55 33.80 13.04 8.12
CodeNER 5.04 10.71 4.19 0.00 3.21 27.52 36.30 34.08 1.59 22.35 28.83 40.15 36.96 9.85 19.80
CodeNER w/o label 8.81 13.47 11.76 2.20 5.21 21.84 31.29 21.05 9.14 19.05 29.00 40.61 37.18 3.98 20.69
Arabic B Vanilla 0.00 0.00 0.00 0.00 0.00 4.62 5.78 5.33 9.68 2.41 33.68 62.17 49.29 14.29 7.35
GoLLIE 15.83 23.67 26.40 6.39 5.58 16.53 27.07 18.98 9.33 8.80 25.12 37.79 34.29 11.65 12.97
GNER 2.14 3.45 1.23 2.74 1.64 2.34 4.62 2.72 0.00 0.89 24.20 39.53 44.36 1.63 5.22
CodeNER 6.79 11.76 2.88 11.43 4.69 27.52 36.30 34.08 1.59 22.35 32.16 46.15 37.25 17.20 21.78
CodeNER w/o label 9.09 13.53 17.07 0.00 2.92 24.69 36.31 31.15 5.85 16.18 33.62 45.52 40.14 15.66 24.45
Finnish A Vanilla 1.92 1.08 0.00 3.95 0.00 23.80 42.90 20.74 27.99 1.71 39.78 56.03 52.20 44.30 0.87
GoLLIE 23.48 28.22 16.89 33.52 3.24 21.86 22.05 17.80 31.99 2.15 34.15 43.23 44.65 38.38 4.07
GNER 18.40 13.64 24.73 23.63 2.20 7.50 9.49 9.88 8.29 1.13 34.80 44.96 40.57 40.71 5.44
CodeNER 27.34 39.63 34.71 27.44 8.28 29.46 31.27 42.03 29.02 14.11 48.89 45.72 70.74 50.94 20.63
CodeNER w/o label 31.69 42.40 44.13 32.61 5.95 32.11 21.01 45.98 39.24 7.58 48.98 43.44 72.01 51.63 19.81
DaNE Vanilla 0.00 0.00 0.00 0.00 0.00 13.49 17.35 9.43 18.09 4.84 13.50 17.82 11.54 20.00 0.00
GoLLIE 27.37 44.23 25.95 26.28 4.85 25.33 38.40 25.20 25.71 5.48 47.17 64.00 56.35 46.87 15.23
GNER 31.42 56.27 14.36 30.86 8.75 26.92 45.89 22.15 27.86 1.26 39.15 44.37 60.82 47.65 2.86
CodeNER 33.79 54.15 53.59 19.72 6.52 37.55 42.29 59.05 36.13 15.33 42.93 50.80 63.33 39.90 19.08
CodeNER w/o label 41.27 61.10 58.33 30.48 12.57 30.02 32.61 49.56 34.11 5.22 35.06 38.22 54.81 33.54 16.74
Average Vanilla 2.72 4.22 2.67 1.92 0.25 26.00 35.78 30.90 26.27 4.75 36.72 52.74 47.67 34.65 8.03
GoLLIE 22.50 32.22 24.76 21.05 3.81 20.74 30.2 22.05 19.67 4.11 34.10 44.50 47.17 29.88 10.17
GNER 21.45 29.87 20.94 18.12 4.48 19.50 27.04 21.25 13.28 1.05 34.59 46.98 46.30 27.44 7.70
CodeNER 22.26 33.99 29.92 14.73 7.44 31.81 37.74 43.38 19.44 12.23 39.17 49.23 52.08 30.94 17.51
CodeNER w/o label 25.64 36.47 35.38 17.98 9.79 31.88 35.45 41.35 22.52 9.73 37.07 45.17 49.53 26.15 17.75
Table 4: Experimental results for each label on FIN, CoNLL03, and SwissNER using the Phi- and Llama-3 models. CodeNER w/o label indicates that we removed entity-label descriptions from the prompt.
5 Analysis
5.1 Open Models
We evaluated the proposed CodeNER with the open models, Phi-3 and Llama-3. We additionally compared our CodeNER with the prompts from zero-shot NER models, GoLLIE and GNER, which are partial code-based prompts that focus on context information and BIO output schema without code, respectively sainz2024gollie; ding-etal-2024-rethinking. Table 4 shows the $F_{1}$ scores, including the detailed performance for each label. For both models, the proposed CodeNER consistently outperformed Vanilla with an average improvement of 19.54, 5.81, and 2.45 points in the $F_{1}$ scores.
Significant differences were observed between the Phi-3 and Llama-3-8B models. We think that the differences in performance between Phi-3 and Llama-3-8B are closely related to the capability of the models to interpret programming language instructions. While Llama-3-8B appears to have a better understanding of the programming language instructions on FIN and Arabic, particularly for NER tasks that require long-range scopes, Phi-3 may require fine-tuning approaches for further improvement. Including entity-label descriptions is not necessary because the averaged scores of CodeNER and CodeNER w/o label exhibit similar performance.
CodeNER significantly outperformed GoLLIE and GNER prompts. By employing a structured code-based approach, CodeNER leverages both BIO output schema and input sentence information, demonstrating that leveraging code is important for improved performance. We also conducted experiments with an increased model size using Llama-3-70B. Our CodeNER outperformed the current prompts from zero-shot NER models of GoLLIE and GNER.
The results presented in Table 10, including prompts with curated descriptions and vanilla outputs, highlight the effectiveness of our CodeNER.
5.2 Impact of BIO Schema
To investigate the effectiveness of incorporating BIO schema within code-based prompts, we evaluated only the “B” and “O” tags, which are crucial for accurately identifying entity span boundaries. Table 5 shows the results. Our CodeNER demonstrates the importance of integrating both the BIO output schema and broader context information of input. CodeNER outperformed the text-based prompt approach of Vanilla. Furthermore, CodeNER surpassed the code-based prompt method of GoLLIE by effectively recognizing start and end tokens rather than relying solely on input sentence information.
Model Dataset Template $F_{1}$ PER LOC ORG MISC O
Phi-3 FIN Vanilla 23.55 4.55 4.88 10.00 0.00 98.33
GoLLIE 25.86 11.81 16.09 9.40 0.00 92.02
GNER 36.66 32.62 32.48 22.30 0.44 95.46
CodeNER 32.71 18.25 34.48 13.10 0.00 97.71
CodeNER w/o label 31.13 13.11 25.81 18.67 0.00 98.05
CoNLL03 Vanilla 27.34 22.86 16.20 6.14 0.51 91.00
GoLLIE 43.54 52.38 37.96 38.10 3.08 86.20
GNER 51.28 72.62 52.07 38.34 7.42 85.96
CodeNER 51.26 58.36 47.97 22.90 33.28 93.78
CodeNER w/o label 58.48 70.34 56.60 36.84 33.63 94.96
SwissNER Arabic A & B Finnish A DaNE Vanilla 19.95 1.31 0.86 3.08 0.34 94.16
GoLLIE 37.00 35.30 30.55 24.55 6.86 87.72
GNER 28.95 16.37 14.90 14.15 5.50 93.85
CodeNER 38.38 38.44 29.63 18.49 10.57 94.76
CodeNER w/o label 41.03 40.08 35.59 22.46 12.23 94.78
Average Vanilla 21.52 4.85 3.63 4.50 0.32 94.30
GoLLIE 36.34 34.38 29.54 24.32 5.34 88.12
GNER 33.96 28.45 24.02 19.54 4.97 92.80
CodeNER 39.41 38.40 32.94 18.35 12.30 95.04
CodeNER w/o label 42.11 40.55 37.19 23.97 13.54 95.27
Llama-3 FIN Vanilla 36.79 19.61 43.04 20.83 3.23 97.22
GoLLIE 27.51 11.84 19.35 10.96 0.00 95.40
GNER 33.66 45.78 14.58 10.33 0.00 97.62
CodeNER 36.91 43.82 34.41 8.33 1.80 96.20
CodeNER w/o label 42.89 59.88 43.40 12.32 3.33 95.55
CoNLL03 Vanilla 65.82 89.09 71.81 56.20 15.08 96.91
GoLLIE 43.28 49.48 40.62 32.63 4.13 89.54
GNER 46.16 52.56 51.45 31.30 1.91 93.56
CodeNER 51.62 54.81 52.80 38.66 17.82 94.02
CodeNER w/o label 49.45 49.77 49.21 40.02 14.91 93.32
SwissNER Arabic A & B Finnish A DaNE Vanilla 35.53 29.25 23.32 25.56 4.72 94.80
GoLLIE 34.87 31.08 23.50 22.48 6.70 90.61
GNER 27.96 15.42 17.16 13.59 1.94 91.69
CodeNER 45.46 41.67 47.31 27.10 16.89 94.32
CodeNER w/o label 42.93 34.89 43.70 29.49 13.44 93.13
Average Vanilla 40.04 36.42 33.06 29.26 5.99 95.45
GoLLIE 35.02 30.96 25.35 22.28 5.38 91.14
GNER 31.94 26.67 22.44 16.00 1.61 92.99
CodeNER 45.12 43.86 46.25 26.07 14.87 94.54
CodeNER w/o label 43.86 40.59 44.45 28.54 12.20 93.50
Table 5: Experimental results for only “B” and “O” tags for each label using the Phi- and Llama-3 models. For SwissNER, Arabic, Finnish, and DaNE, we averaged their scores.
Figure 3: Case studies of Vanilla and CodeNER outputs. Text denotes the input sentence, and Gold refers to the correct label for Text. Vanilla Output denotes the results generated by Vanilla. Vanilla Result represents the conversion of Vanilla Output into the BIO-tagging format for evaluation. CodeNER Result denotes the CodeNER labeling results for Text.
5.3 Case Study
We conducted a case study of the outputs from CodeNER and Vanilla. The results are shown in Figure 3. In Figure 3 (a), Vanilla failed to treat the website as a single token; instead, it extracted and labeled individual words. In contrast, the proposed CodeNER accurately recognized the website as a single token. In Figure 3 (b), while Vanilla generated duplicate labels for the same word, CodeNER did not suffer from this problem because it can capture the structured token-by-token approach. In Figure 3 (c), when a word is repeated multiple times in the text, Vanilla tended to annotate labels only once, whereas CodeNER did not. In Figure 3 (d), Vanilla could not capture special tokens, such as comma and period, when attached to a word because it relies solely on sentence-level information. However, CodeNER successfully identified special tokens. Details of evaluations for such cases are described in Appendix B.
5.4 Combination with Chain-of-thought
We evaluated the proposed CodeNER by incorporating the CoT reasoning instruction, “Let’s think step by step”, at the beginning of the CodeNER instruction. Although previous text-based vanilla prompting methods have been also explored to incorporate CoT in NER han2023informationextractionsolvedchatgpt; xie-etal-2023-empirical, the proposed approach, CodeNER, presents a different perspective. It encourages LLMs to enhance their understanding of long-range scopes using programming language-style prompts. Table 6 shows the results. The integration of CoT with CodeNER demonstrated the potential to improve performance in zero-shot settings by leveraging structured programming language-style prompts.
5.5 Different Programming Languages
It is also of interest to perform a comparison of different programming languages such as C++. [^7] CodeNER with C++ prompts is in Appendix F. Table 7 shows the results. CodeNER with C++ instructions showed significant improvements in the $F_{1}$ scores over both Vanilla and CodeNER with Python instructions on FIN. Although the consideration of C++ showed better person entity recognition performance on FIN, the differences between the two programming languages in prompts were less noticeable.
Interestingly, the $F_{1}$ scores with CodeNER remain consistent between the Python and C++ instructions on CoNLL03. We think this suggests that dataset characteristics may have a more significant impact than the choice of the programming language. The comparable results between Python and C++ suggest that the choice of the programming language has a minimal impact on the effectiveness of CodeNER when used on datasets where the model’s existing capabilities are strong. We additionally considered using the Java programming language for CodeNER. However, the performance was nearly zero in terms of the F1 scores, which indicates LLMs lack inherent knowledge of Java in their pre-training haroon2025accuratelylargelanguagemodels.
Model Dataset Instruction $F_{1}$ PER LOC ORG MISC
GPT-4 FIN Vanilla 22.83 26.94 15.62 14.72 1.06
Vanilla+CoT 24.29 29.27 11.94 16.59 0.91
GoLLIE 12.59 8.75 44.16 7.14 1.91
GoLLIE+CoT 12.52 8.93 42.67 7.21 0.89
CodeNER 41.13 51.71 38.10 7.48 0.78
CodeNER+CoT 42.09 52.69 38.60 8.80 0.87
CoNLL03 Vanilla 73.24 92.06 79.45 66.15 31.88
Vanilla+CoT 72.53 91.71 79.61 64.64 30.18
GoLLIE 51.96 65.97 62.48 46.89 6.67
GoLLIE+CoT 51.76 65.62 62.22 46.85 6.57
CodeNER 69.48 88.38 75.77 58.53 36.87
CodeNER+CoT 69.32 87.57 76.47 57.86 37.38
Phi-3 FIN Vanilla 1.24 1.82 0.00 0.00 0.00
Vanilla+CoT 1.84 2.70 0.00 0.00 0.00
GoLLIE 8.26 9.59 6.90 5.02 0.00
GoLLIE+CoT 4.66 4.27 7.23 4.82 1.11
CodeNER 17.51 18.18 27.12 10.43 0.00
CodeNER+CoT 18.73 19.61 28.57 10.86 0.00
CoNLL03 Vanilla 11.73 21.20 15.44 3.75 0.00
Vanilla+CoT 12.75 22.27 16.43 4.96 0.53
GoLLIE 35.39 51.47 33.88 35.43 1.84
GoLLIE+CoT 27.13 37.23 27.37 27.68 1.97
CodeNER 36.57 54.03 44.64 19.75 16.95
CodeNER+CoT 44.62 61.94 54.63 26.98 22.66
Llama-3 FIN Vanilla 18.14 16.47 35.44 15.00 0.00
Vanilla+CoT 18.47 16.34 37.50 15.77 1.56
GoLLIE 9.21 9.87 8.60 8.22 0.00
GoLLIE+CoT 9.92 10.29 14.29 6.14 4.49
CodeNER 31.96 39.22 34.04 6.54 0.00
CodeNER+CoT 27.79 35.15 22.22 6.77 0.00
CoNLL03 Vanilla 64.02 88.90 71.01 54.85 11.85
Vanilla+CoT 64.08 88.70 70.70 54.52 14.26
GoLLIE 34.72 49.03 38.32 30.65 2.84
GoLLIE+CoT 38.53 52.08 43.78 35.23 2.67
CodeNER 40.29 51.07 50.70 31.01 12.71
CodeNER+CoT 44.21 55.68 56.68 32.32 16.29
Table 6: Experimental results including each label on FIN and CoNLL03 by incorporating CoT.
Model Dataset Instruction $F_{1}$ PER LOC ORG MISC
GPT-4 FIN Vanilla 22.83 26.94 15.62 14.72 1.06
GoLLIE 12.59 8.75 44.16 7.14 1.91
CodeNER (Python) 41.13 51.71 38.10 7.48 0.78
CodeNER (C++) 44.44 56.84 33.33 9.85 0.55
CoNLL03 Vanilla 73.24 92.06 79.45 66.15 31.88
GoLLIE 51.96 65.97 62.48 46.89 6.67
CodeNER (Python) 69.48 88.38 75.77 58.53 36.87
CodeNER (C++) 68.43 85.98 74.44 58.59 37.01
Phi-3 FIN Vanilla 1.24 1.82 0.00 0.00 0.00
GoLLIE 8.26 9.59 6.90 5.02 0.00
CodeNER (Python) 17.51 18.18 27.12 10.43 0.00
CodeNER (C++) 14.61 14.05 12.00 20.41 0.00
CoNLL03 Vanilla 11.73 21.20 15.44 3.75 0.00
GoLLIE 35.39 51.47 33.88 35.43 1.84
CodeNER (Python) 36.57 54.03 44.64 19.75 16.95
CodeNER (C++) 38.44 56.84 47.63 21.57 14.15
Llama-3 FIN Vanilla 18.14 16.47 35.44 15.00 0.00
GoLLIE 9.21 9.87 8.60 8.22 0.00
CodeNER (Python) 31.96 39.22 34.04 6.54 0.00
CodeNER (C++) 25.08 31.62 19.23 7.07 0.00
CoNLL03 Vanilla 64.02 88.90 71.01 54.85 11.85
GoLLIE 34.72 49.03 38.32 30.65 2.84
CodeNER (Python) 40.29 51.07 50.70 31.01 12.71
CodeNER (C++) 32.84 42.23 44.05 20.28 14.25
Table 7: Experimental results including each label on FIN and CoNLL03 with different programming languages.
5.6 Trade-off Discussion
Advantages. The proposed CodeNER provides several advantages because it successfully bridges the gap between the text-in-text-out schema of LLMs and the text-in-span-out nature for handling the sequential aspect of BIO output schema. The Vanilla prompt method, during labeling, leads to overlapping labels, which results in unclear classifications; in contrast, CodeNER leverages sequential token information through a for-loop, which avoids errors from overlapping labels and ensures that each token is labeled only once. In cases requiring the identification of individual tokens, CodeNER accurately recognizes a token as a single unit, whereas the Vanilla prompt method struggles to label tokens such as web addresses. Therefore, for datasets requiring NER to recognize single labels, CodeNER is more effective. In addition, when the same word appears multiple times, the Vanilla prompt faces difficulties in determining which tokens to label. However, CodeNER accurately labels each token.
Disadvantages. Although CodeNER effectively recognizes named entities, there are cases in which considering the BIO output schema within code blocks may not be advantageous. In our in-depth analysis of two datasets, MIT-Movie and MIT-restaurant, we found that MIT-Movie contains long sentences in which the labels “Plot” and “Quote” appear frequently. These labels occupy significant portions (42.61%) of the dataset. They typically encompass entire sentences or long text spans that include function words such as conjunctions, prepositions, and determiners, for example, “in,” “is,” “with,” and “the”, while such words do not convey specific meanings. Consequently, for the datasets where sentences or long text spans with function words need to be labeled, the Vanilla prompt method, which considers only input sentence information, can be more effective than CodeNER. [^8] Additional case studies are described in Appendix G.
6 Conclusion
In this study, we introduced CodeNER, a novel code-based prompting method designed to enhance the capabilities of LLMs in performing NER. By embedding structured code within prompts, CodeNER effectively integrates the BIO output schema and input sentence information, thereby leveraging the inherent strengths of programming language comprehension in LLMs. Experimental results on ten NER benchmarks demonstrated that CodeNER consistently outperforms text-based prompting methods by bridging the inherent gap between NER and LLMs.
Limitations
Although CodeNER showed significant improvements by leveraging code-based prompting for NER, it relies on the capabilities of LLMs. LLMs without being trained in programming languages may require further fine-tuning. Moreover, for datasets containing long sentences or long text spans with function words, the Vanilla prompt method that considers only input sentence information can sometimes be more effective than our CodeNER. In the future, we plan to pre-train our CodeNER by incorporating a related auxiliary task that focuses on identifying function words or syntactic structures across various datasets.
Appendix A Hyperparameters
Table 8 shows the hyperparameters and computing settings used in our experiments.
Appendix B Evaluation Details
When Vanilla generated duplicate outputs, we considered the longest one for evaluating the F1 score (case (a) in Figure 3). In addition, when the Vanilla generated output tags only once despite there being multiple gold outputs (case (c) in Figure 3), we considered the first one for evaluation. Note that there was no significant difference between considering the shortest or the last outputs for evaluation in that they produce comparable results.
Llama-3 (8B) Phi-3 (mini-4k-instruct)
Computing Setting NVIDIA RTX A6000 NVIDIA RTX A6000
Quantization 8-bit 8-bit
Batch size 1 1
do_sample False False
Max_new_tokens 1200 1200
GPT-4 GPT-4 Turbo
max_tokens 1500 1500
Temperature 0 0
Table 8: Hyperparameters in the experiments.
Appendix C GoLLIE Prompt
Figure 4: Examples of prompts for Vanilla with label descriptions. (a) is the examples of Vanilla with label same as codeNER and (b) is the examples of Vanilla with label generated by GPT.
Figure 5: GoLLIE prompt for CoNLL03, which includes labels such as PER, ORG, LOC, and MISC.
Figure 5 shows GoLLIE prompt.
Appendix D Result on Few-shot settings
We provide experimental results on one-shot and three-shot settings for models, Phi-3-mini-128k-instruct and lama3-8B-instruct. Table 9 shows the results. For one-shot and three-shot settings, we randomly selected instances from each train dataset.
Appendix E Result on Vanilla w/ label
We conducted additional experiments by augmenting the Vanilla model with curated label descriptions. Specifically, we compared two descriptions: one generated by GPT-o3-mini-high (shown in Figure 4) and another is used for our CodeNER method. Table 10 reports the results. Even when Vanilla considered carefully crafted descriptions, incorporating CodeNER yields a clear performance gain.
Appendix F CodeNER with C++
Figure 6: CodeNER using C++.
Figure 6 presents CodeNER with the C++ instruction.
Appendix G Additional Case study
Figure 7: Case of labeling function word in Vanilla and CodeNER
We conducted a detailed case study of the outputs of Vanilla and CodeNER for the “Plot” label on MIT-Movie. In Figure 7, the proposed CodeNER faces difficulties in handling determiners and prepositions, such as “the” and “in” due to reinforcing the token-level information.
Setting Model Dataset Template $F_{1}$ PER LOC ORG MISC Model $F_{1}$ PER LOC ORG MISC
One-shot Phi-3 CoNLL03 Vanilla 57.57 81.84 62.67 52.24 2.18 Llama-3 8B 56.90 84.56 60.25 46.45 9.95
CodeNER 52.90 71.38 59.12 44.42 15.62 48.04 64.64 51.48 40.44 19.65
FIN Vanilla 28.21 37.42 7.06 10.39 0.00 28.53 35.04 15.62 15.64 0.00
CodeNER 43.09 53.44 33.85 14.29 1.65 49.87 63.18 46.67 6.10 2.34
Arabic A Vanilla 4.06 11.43 2.97 0.00 1.27 21.59 17.58 9.23 12.50 11.68
CodeNER 19.99 27.68 22.79 0.00 16.80 28.87 44.53 35.64 10.11 18.29
Arabic B Vanilla 6.94 5.46 2.63 0.00 11.88 15.00 34.68 15.46 5.06 2.36
CodeNER 14.02 15.00 11.69 10.17 15.42 29.79 29.79 43.48 8.40 18.59
Finnish A Vanilla 25.42 36.31 9.03 39.78 1.33 23.99 50.45 10.76 28.99 5.97
CodeNER 40.36 49.88 52.57 45.05 6.80 34.47 40.10 56.64 31.03 11.83
DaNE Vanilla 30.19 52.17 41.23 20.34 1.85 54.92 76.32 68.42 61.18 4.03
CodeNER 45.26 72.51 66.01 24.47 15.92 40.60 53.02 65.18 33.71 11.81
Average Vanilla 25.40 37.44 20.93 20.46 3.09 33.49 49.77 29.96 28.30 5.67
CodeNER 35.94 48.32 41.01 23.07 12.04 38.61 50.82 49.85 21.63 13.75
Three-shot Phi-3 CoNLL03 Vanilla 59.36 78.55 68.29 55.70 2.58 Llama-3 8B 60.31 86.87 70.46 45.40 10.30
CodeNER 52.26 73.12 55.99 40.61 22.92 47.70 57.06 53.42 43.34 22.85
FIN Vanilla 29.24 35.53 24.24 11.76 0.40 30.44 36.30 24.56 15.47 0.00
CodeNER 39.08 53.44 33.85 14.29 1.65 47.90 59.38 49.35 7.74 3.08
Arabic A Vanilla 6.23 13.95 5.00 3.20 3.16 19.27 49.76 5.21 13.19 12.54
CodeNER 23.03 30.25 30.60 2.13 17.00 29.84 44.79 37.59 0.00 20.69
Arabic B Vanilla 4.98 8.21 4.97 5.00 2.64 10.17 12.79 20.95 10.45 1.76
CodeNER 22.86 33.44 27.78 5.94 15.42 31.51 41.58 34.68 4.35 27.39
Finnish A Vanilla 26.10 38.22 14.33 38.14 1.16 37.90 51.64 41.40 43.91 8.06
CodeNER 43.83 54.09 51.32 50.32 10.96 40.17 39.02 57.77 42.51 14.63
DaNE Vanilla 35.77 56.30 41.55 33.45 3.72 48.73 71.98 46.88 52.07 11.18
CodeNER 45.62 69.66 62.14 33.60 12.77 42.53 49.90 60.83 43.10 16.29
Average Vanilla 26.95 38.46 26.40 24.54 2.28 34.47 51.56 34.91 30.08 7.31
CodeNER 37.78 52.33 43.61 24.48 13.45 39.94 48.62 48.94 23.51 17.49
Table 9: Experimental results of one-shot and three-shot for each label on FIN, CoNLL03, Arabic A B, Finnish A, and DaNE using the Phi- and Llama-3 models.
Model Dataset Template $F_{1}$ PER LOC ORG MISC Model $F_{1}$ PER LOC ORG MISC
Phi-3 FIN Vanilla 1.24 1.82 0.00 0.00 0.00 Llama-3 8B 18.14 16.47 35.44 15.00 0.00
Vanilla w/ label (GPT) 1.85 2.71 0.00 0.00 0.00 18.31 16.80 32.94 16.43 0.00
Vanilla w/ label (codeNER) 1.24 1.82 0.00 0.00 0.00 18.40 16.47 36.36 15.64 1.75
CodeNER 17.51 18.18 27.12 10.43 0.00 31.96 39.22 34.04 6.54 0.00
CodeNER w/o label 8.49 7.35 19.05 6.62 0.00 45.62 57.26 41.12 9.40 1.33
CoNLL03 Vanilla 11.73 21.20 15.44 3.75 0.00 64.02 88.90 71.01 54.85 11.85
Vanilla w/ label (GPT) 11.92 21.50 15.33 3.86 0.83 64.73 89.32 72.32 52.95 17.91
Vanilla w/ label (codeNER) 9.38 15.49 9.12 7.55 0.24 64.24 89.13 71.25 53.89 14.76
CodeNER 36.57 54.03 44.64 19.75 16.95 40.29 51.07 50.70 31.01 12.71
CodeNER w/o label 48.33 66.92 55.66 32.26 26.12 37.00 44.27 46.17 31.80 10.75
SwissNER Vanilla 4.14 5.45 3.23 5.71 1.72 49.04 72.91 58.82 52.06 6.15
Vanilla w/ label (GPT) 1.98 3.74 1.09 3.02 0.00 52.28 73.06 67.23 52.91 8.00
Vanilla w/ label (codeNER) 1.63 7.41 0.00 1.02 0.00 56.54 70.83 70.83 61.58 11.76
CodeNER 28.76 49.47 42.34 14.35 12.44 35.48 28.99 62.37 30.55 6.55
CodeNER w/o label 31.78 50.55 41.67 21.71 15.76 31.87 25.37 54.41 28.09 8.02
Arabic A Vanilla 0.00 0.00 0.00 0.00 0.00 8.88 6.13 15.53 6.25 6.31
Vanilla w/ label (GPT) 0.00 0.00 0.00 0.00 0.00 10.53 18.18 14.08 13.21 3.38
Vanilla w/ label (codeNER) 0.00 0.00 0.00 0.00 0.00 16.80 25.40 20.10 7.55 11.11
CodeNER 5.04 10.71 4.19 0.00 3.21 27.52 36.30 34.08 1.59 22.35
CodeNER w/o label 8.81 13.47 11.76 2.20 5.21 21.84 31.29 21.05 9.14 19.05
Arabic B Vanilla 0.00 0.00 0.00 0.00 0.00 4.62 5.78 5.33 9.68 2.41
Vanilla w/ label (GPT) 0.00 0.00 0.00 0.00 0.00 9.87 11.70 14.01 17.72 4.60
Vanilla w/ label (codeNER) 0.00 0.00 0.00 0.00 0.00 10.92 20.51 6.76 6.98 7.22
CodeNER 6.79 11.76 2.88 11.43 4.69 27.52 36.30 34.08 1.59 22.35
CodeNER w/o label 9.09 13.53 17.07 0.00 2.92 24.69 36.31 31.15 5.85 16.18
Finnish A Vanilla 1.92 1.08 0.00 3.95 0.00 23.80 42.90 20.74 27.99 1.71
Vanilla w/ label (GPT) 15.95 7.80 8.52 29.01 0.00 25.91 46.21 13.77 34.50 3.00
Vanilla w/ label (codeNER) 13.35 7.88 2.79 25.97 0.00 27.15 48.56 18.40 33.94 3.65
CodeNER 27.34 39.63 34.71 27.44 8.28 29.46 31.27 42.03 29.02 14.11
CodeNER w/o label 31.69 42.40 44.13 32.61 5.95 32.11 21.01 45.98 39.24 7.58
DaNE Vanilla 0.00 0.00 0.00 0.00 0.00 13.49 17.35 9.43 18.09 4.84
Vanilla w/ label (GPT) 0.00 0.00 0.00 0.00 0.00 21.80 22.79 30.22 29.86 2.92
Vanilla w/ label (codeNER) 0.00 0.00 0.00 0.00 0.00 8.44 9.80 14.95 4.71 6.20
CodeNER 33.79 54.15 53.59 19.72 6.52 37.55 42.29 59.05 36.13 15.33
CodeNER w/o label 41.27 61.10 58.33 30.48 12.57 30.02 32.61 49.56 34.11 5.22
Average Vanilla 2.72 4.22 2.67 1.92 0.25 26.00 35.78 30.90 26.27 4.75
Vanilla w/ label (GPT) 4.53 5.11 3.56 5.13 0.12 29.06 39.72 34.94 31.08 5.69
Vanilla w/ label (codeNER) 3.66 4.66 1.70 4.93 0.03 28.93 40.10 34.09 26.33 8.06
CodeNER 22.26 33.99 29.92 14.73 7.44 31.81 37.74 43.38 19.44 12.23
CodeNER w/o label 25.64 36.47 35.38 17.98 9.79 31.88 35.45 41.35 22.52 9.73
Table 10: Experimental results for each label on FIN, CoNLL03, and SwissNER using the Phi- and Llama-3 models. Vanilla w/ label (GPT) and Vanilla w/ label (codeNER) each indicate that we added the corresponding labels explanation to the Vanilla prompt, labels explanation generated by GPT for the former, and labels explanation provided by CodeNER for the latter.